OpenStudy (anonymous):

solve the equation show your work 27^x ^- 6 = 9^x^ – 1

OpenStudy (whpalmer4):

Can you rewrite that with the equation editor? Click the \(\sum\) Equation button at lower left...

OpenStudy (anonymous):

it only shows up like tht...but x-6 is sopose to be to the power like an exponet as well as x-1

OpenStudy (whpalmer4):

Oh, so it is this? \[27^{x-6} = 9^{x-1}\] The way you've written suggests you meant \[27^{x^{-6}} = 9^{x^{-1}}\]

OpenStudy (anonymous):

oh sorry its the 1st one you wrote

OpenStudy (whpalmer4):

the answer is much nicer for the 1st option :-)

OpenStudy (anonymous):

yea so can u help?

OpenStudy (xanthe):

make the base the same...

OpenStudy (whpalmer4):

Why don't you divide both sides by \(9^{x-1}\)? You can use the property of exponents that \[(ab)^n = a^nb^n\] to split the left side into a 3^ part * a 9^ part

OpenStudy (whpalmer4):

Then remember that \(a^0 = 1\) so long as \(a \ne 0\).

OpenStudy (xanthe):

You do the rest